Cede & Co. expects its EBIT to be $64,000 every year forever. The firm can borrow at 8 percent. The firm currently has no debt, its cost of equity is 14 percent, and the tax rate is 35 percent. Assume the firm borrows $171,000 and uses the proceeds to repurchase shares. What is the cost of equity after recapitalization? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) Cost of equity % What is the WACC? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) WACC %
Answer #1

Value of Unlevered Firm = EBIT * (1 - Tax Rate) / Unlevered Cost of Equity
Value of Unlevered Firm = $64,000 * (1 - 0.35) / 0.14
Value of Unlevered Firm = $41,600 / 0.14
Value of Unlevered Firm = $297,142.86

Value of Levered Firm = Value of Unlevered Firm + Tax Rate * Value of Debt
Value of Levered Firm = $297,142.86 + 0.35 * $171,000
Value of Levered Firm = $297,142.86 + $59,850
Value of Levered Firm = $356,992.86

Value of Equity = Value of Levered Firm - Value of Debt
Value of Equity = $356,992.86 - $171,000
Value of Equity = $185,992.86

Weight of Debt = Value of Debt / Value of Levered Firm
Weight of Debt = $171,000 / $356,992.86
Weight of Debt = 0.4790

Weight of Equity = Value of Equity / Value of Levered Firm
Weight of Equity = $185,992.86 / $356,992.86
Weight of Equity = 0.5210

Debt-Equity Ratio = Value of Debt / Value of Equity
Debt-Equity Ratio = $171,000 / $185,992.86
Debt-Equity Ratio = 0.91939

Levered Cost of Equity = Unlevered Cost of Equity + (Unlevered Cost of Equity - Cost of Debt) * (1 - Tax Rate) * Debt-Equity Ratio
Levered Cost of Equity = 0.14 + (0.14 - 0.08) * (1 - 0.35) * 0.91939
Levered Cost of Equity = 0.14 + 0.0359
Levered Cost of Equity = 0.1759 or 17.59%

WACC = Weight of Debt * Cost of Debt * (1 - Tax Rate) + Weight of Equity * Cost of Equity
WACC = 0.4790 * 8.00% * (1 - 0.35) + 0.5210 * 17.59%
WACC = 11.66%